the sticker on nestle's crunch bar reads 20.4 grammes. Suppose that the probability distribution of the weight crunch bar is known to follow a normal distribution with variance of 0.16 and mean 21.37 grammes. Calculatethe probability that a weight of a crunch bar selected from the shelf will exceed 22.07 grammes


1
Expert's answer
2021-09-02T23:08:33-0400

Solution:

"X\sim N(\mu,\sigma^2)
\\\mu=21.37,\sigma^2=0.16"

"P(X>22.07)=1-P(X\le22.07)
\\=1-P(z\le \dfrac{22.07-21.37}{\sqrt{0.16}})
\\=1-P(z\le1.75)
\\=1-0.95994
\\=0.04006"